Find an equation of the plane passing through the given points. (3, 4, -4), (3, -4,4), (-3, -4, -4) 7x - 6y - 7z - 25 = 0x
Answer #1

Equation of a plane passing through points 21,41, 21 ,12. Y2, 22 and 13, 93, 23 is given by following equation containing determinant

1 - 11 12 – Ii 23 – 21 Y - Y1 Y2 – Yi Y3 — Yi 2 – 21 22 – 21 = 0 23 – 21

here points are given . Use this points to find equation of plane

9-4 C-3 3-3 1-3-3 -4-4 -4-4 244 4+4|= 0 -4+4

Y-4 C – 3 0 -6 2 +4 8 0 -8 -8 = 0

now expand through first row to find determinant

(x-3)[0+64]-(y-4)[0+48]+(z+4)[0-48]=0

64x-192-48y+192-48z-192=0

64x-48y-48z-192=0

divide both side by 16 to simplify this . Then equation becomes.

4x-3y-3z-12=0

so this is answer